Hello, could someone please help me solve this problem. Thank you in advance.

Problem:
The perimeter of a rectangle is 36 m. The length is 2 m more than three times the width. Find the length and width of the rectangle.

Found 2 solutions by Horlaye, Alan3354:
Answer by Horlaye(27)

Formula for perimeter of a rectangle is 2(l+w).
From the first sentence
2(l+w) = 36
2l + 2w = 36.........(1)
From the second sentence
l is 2 more than 3w
l - 3w = 2.........(2)
Multiply equation (1) by 3 and (2) by 2
6l + 6w = 108......(4)
2l - 6w = 4.......(5)
Add up both equation
8l = 112
Divide both sides by 8
l = 112/8
l = 14
Substitute l = 14 in equation (2)
l - 3w = 2
14 - 3w = 2
Collect like terms
-3w = 2 - 14
-3w = -12
Divide both sides by -3
w = -12/-3
w = 4
The length and width of the rectangle is 14 and 4 respectively.

Answer by Alan3354(69443)

The perimeter of a rectangle is 36 m. The length is 2 m more than three times the width. Find the length and width of the rectangle.
----------
Perimeter = 2L + 2W = 36
L + W = 18
L = 3W + 2
---
Sub for L in the 1st eqn
(3W + 2) + W = 18
Solve for W
then solve for L